OpenStudy (anonymous):

find the x-intercept of the parabola with vertex (4,-1) and y- intercept (0,15). write answer in (x1,y1) (x2,y2) form :D

OpenStudy (anonymous):

vertex (4,-1) y =a(x-4)^2 -1 using y-intercept pt(0,15) to find a 15 = a(-4)^2 - 1 16 = a*16 1 = a y = (x-4)^2 - 1 OR y = x^2 - 8x + 15 x^2 - 8x + 15 = 0 factoring (x-3)(x-5) = 0 (x-3) = 0 x = 3 (x-5) = 0 x = 5 x-intercepts are Pt(3,0) and pt(5,0)

OpenStudy (anonymous):

do i put (3,0),(5,0) or (3,0) (5,0) ?

OpenStudy (anonymous):

your answer is both (3,0) and (5,0)

OpenStudy (anonymous):

WOOOOOOOOOOOOOOO it was correct

OpenStudy (anonymous):

oh good you had me worried there for a sec

OpenStudy (anonymous):

lol :) thanks so muchhh <33 but i have another one now :( u dont needa do the entire work; i just need the answeer :)

OpenStudy (anonymous):

i was giving you the work, so you could do the next one mostly yourself its not like you're dumb or anything

OpenStudy (anonymous):

the thing is that i am :( thats why im in summer school lol. well can u help me? like give me directtions?

OpenStudy (anonymous):

yes but please dont rush me i have to do other work too

OpenStudy (anonymous):

no rushing ^_^ <3 just tell me here when u can help me.

OpenStudy (anonymous):

its probably better to post the question anyway, cause theres a fair few people on here that could help you ... not just 12 year olds :) i've been getting help with my calculus/matrices/ stuff you learn in college

OpenStudy (anonymous):

i did post it but no one is helping me :( & omg thats cool. thanks for the help anyways.